# Local consistency and axioms of functional dependence

### v1

@@ 3.1 $K$-realisability of Contextual Families
$>$ #$\#(\texttt{Student}\mapsto\text{Bob}),$
where the inequality follows since each value of ##\## is strictly positive. Thus the contextual family of Fig. 1 cannot be obtained as the support of any contextual $\mathbb{R}_{\geq 0}$-family.
Next, we completely characterise which contextual families are supports of some contextual $K$-families for cancellative monoids $K$ and context sets that form a chordless cycle (defined below). We show that a contextual family $\mathcal{R}$ is the support of some contextual $K$-family if and only if its associated overlap projection graph has an edge cycle cover; both terms are defined below.
For what follows, we treat addition and subtraction of natural numbers $i\in[n]$ modulo $n$. We call a context set $\mathcal{C}=\{C_{0},\dots,C_{n-1}\}$, where $n\geq 3$, a chordless-cycle context set if for all $i,j\in[n]$, the contexts $C_{i}$ and $C_{j}$ intersect if and only if $i=j\pm 1$. For the remainder of this section, if the context $\mathcal{C}=\{C_{0},\dots,C_{n-1}\}$ is implicitly understood, we write $\mathcal{S}_{i}$ as a shorthand for $\mathcal{S}_{C_{i}}$, given a contextual $K$-family $\mathcal{S}$ over $\mathcal{C}$.
Let $\mathcal{R}$ be a contextual family over a chordless-cycle context set $\mathcal{C}=\{C_{0},\dots,C_{n-1}\}$. The overlap projection graph of $\mathcal{R}$, denoted $\mathrm{OPGraph}(\mathcal{R})$, is the digraph s.t.:
  • The vertices consists of the restrictions $s\!\restriction\!(C_{i}\cap C_{i+1})$, for each $s\in\mathcal{R}_{i}$, $i\in[n]$, and
@@
A digraph $G=(V,E)$ is called cyclically $n$-partite if there exists a partition $\{V_{0},\dots,V_{n-1}\}$ of $V$ such that for every edge $(u,v)\in E$ we have $u\in V_{i}$ and $v\in V_{i+1}$. The following lemma is a simple observation.
@@
Let $\mathcal{R}$ be a contextual family over a chordless-cycle context set $\mathcal{C}=\{C_{0},\dots,C_{n-1}\}$. Then $\mathrm{OPGraph}(\mathcal{R})$ is cyclically $n$-partite.
A digraph $G$ has an edge cycle cover if every edge $(u,v)\in E$ belongs to some cycle. We will characterise $K$-realisable contextual families with respect to chordless cycles using this notion. The next example showcases the concepts introduced so far.
Example 7.
Consider again the contextual family in Figure 1. The set $\mathcal{C}=\{C_{1},C_{2},C_{3}\}$, where $C_{1}=\{\texttt{Student},\texttt{Teacher}\}$, $C_{2}=\{\texttt{Teacher},\texttt{Course}\}$, and $C_{3}=\{\texttt{Course},\texttt{Student}\}$, is a chordless-cycle context set. The corresponding overlap projection graph, depicted in Figure 2, has $6$ vertices (one for each value) and $7$ edges (one for each assignment). In the figure, we represent a vertex, that is, a restriction $s\upharpoonright\{x\}$, simply as the unique value $s(x)$ it corresponds to. We note that the graph is cyclically $3$-partite, as we can partition the vertices to students, teachers, and courses with edges between them. Moreover, the vertices Alice, Charlie, and Math form a cycle, and similarly Bob, David, and CS form a cycle. Each edge belongs to some cycle, except for the edge from CS to Alice. The graph thus has no edge cycle cover.
MathAliceCharlieCSBobDavid
Figure 2: Example of an Overlap Projection Graph
To characterise $K$-realisable contextual families over chordless cycles, we need the following lemma, which states that every simple cycle gives rise to a uniformly weighted $K$-relation. Recall that a cycle $(v_{1},\dots,v_{\ell},v_{1})$ in a graph is called simple if does not visit any vertex twice. Let us then call a contextual family $\mathcal{R}$ over a chordless-cycle context set simply cyclic if $\mathrm{OPGraph}(\mathcal{R})$ contains a simple cycle
@@
where =$\bigcup{\mathcal{R}}=\{s_{1},\dots,s_{\ell}\}$. The next lemma states that any simply cyclic family can be realised as a $K$-family using uniform multiplicities. {restatable}lemmahelp Let $\mathcal{R}$ be a simply cyclic contextual family over a chordless-cycle context set $\mathcal{C}=\{C_{0},\dots,C_{n-1}\}$. Let $K$ be a positive commutative monoid and $0\neq w\in K$. Then $w\mathcal{R}$ is a contextual $K$-family, and $\mathrm{Supp}(w\mathcal{R})=\mathcal{R}$.
@@
We need to prove that $w\mathcal{R}$ is locally consistent. Suppose the simple cycle has length $\ell$, as in (1). Since $\mathrm{OPGraph}(\mathcal{R})$ is cyclically $n$-partite (Lemma 6), $\ell$ must be a multiple of $n$; that is, $n\cdot m=\ell$ for some positive integer $m$. Fix $i\in[n]$. One observes that
@@
Theorem 8.
Let $K$ be a positive commutative cancellative monoid and $\mathcal{R}$ a contextual family over a chordless-cycle context set $\mathcal{C}=\{C_{0},\dots,C_{n-1}\}$. Then, $\mathcal{R}$ is the support of some contextual $K$-family if and only if its overlap projection graph has an edge cycle cover.
@@
( $\Rightarrow$ ) Let $\mathcal{R}$ be the support of some contextual $K$-family $\mathcal{S}$. Then $\mathcal{S}$ can be represented as #$\#\colon\bigcup{\mathcal{R}}\to K_{\neq 0}$, a mapping which we extend to sets $S$ by #$\#(S)\coloneqq\sum_{s\in S}\#(s)$. Toward a contradiction, assume there is no edge cycle cover for $\mathrm{OPGraph}(\mathcal{R})$. Then, some $s\in\bigcup{\mathcal{R}}$ gives rise to an edge that is not in any cycle. Without loss of generality, we assume that $s$ is associated with the context $C_{0}$, that is, it belongs to $\mathcal{R}_{0}$. We define sets of assignments $L^{j}$, $j\geq 0$, inductively as follows:
  • $L_{0}\coloneqq\{s^{\prime}\in\mathcal{R}_{{1}}\mid{s}\!\restriction\!{(C_{0}\cap C_{1})}={s^{\prime}}\!\restriction\!{(C_{0}\cap C_{1})}\}$,
  • $L^{j+1}$ extends $L^{j}$ with assignments $s^{\prime}\in\mathcal{R}_{{j+1}}$, $j\in[n]$, such that there is $s\in L_{i}\cap\mathcal{R}_{{j}}$ with
@@
#$\#(M_{0}\cup\{s\})\leq\#(M_{1})\leq\ldots\leq\#(M_{n-1})\leq\#(M_{0}),$ (2)
where $\leq$ refers to the natural (pre)order of $K$. Note that if this claim holds, we reach a contradiction. Indeed, note that $s\notin L_{0}$ by hypothesis, hence the transitivity of $\leq$ implies #$\#(M_{0})+\#(\{s\})=\#(M_{0}\cup\{s\})\leq\#(M_{0})$. Thus, there is $c\in K$ such that #$\#(M_{0})+\#(\{s\})+c=\#(M_{0})$. By cancellativity, #($\#(\{s\})+c=0$, and hence by positivity, #$\#(s)=0$. This contradicts the fact that $s\in\bigcup{\mathcal{R}}$, where ##\## is a function from $\bigcup{\mathcal{R}}$ to $K_{\neq 0}$.

### v2

@@ 4.2 Contextual $\mathbb{R}_{\geq 0}$-realisability
Figure 4: Non-realisable contextual family.
Proposition 20.
Proof.
Corollary 21.
An axiom system for FDs is sound (complete, resp.) with respect to $\models_{\mathbb{R}_{\geq 0}}$ if and only it is that for $\models_{\mathbb{N}}$.
5 A Path to More General Axiomatisations
We have shown a sound and complete axiomatisation for FDs restricted to unary FDs and binary CDs. We also showed that the system consisting of reflexivity, augmentation, and contextual transitivity is the strongest uni-context axiomatisation. Next we scout further avenues for more general axiomatisations, and present some preliminary results.
5.1 Unary FDs with Ternary Contexts
For unary FDs with ternary CDs, the interaction is significantly more intricate. We have already observed that transitivity can fail: from $x_{1}\rightarrow x_{2}$ and $x_{2}\rightarrow x_{3}$, it does not necessarily follow that $x_{1}\rightarrow x_{3}$. However, if we additionally assume $c\rightarrow x_{3}$, $cx_{1}x_{2}$, $cx_{2}x_{3}$, and $cx_{1}x_{3}$, then $x_{1}\rightarrow x_{3}$ does follow. The next rule captures this non-trivial interaction between unary FDs and ternary CDs (the above case is its instance with $n=3$ and $c_{i}=c$ for all $i\in[n]$).
$c_{n-2}x_{n-1}c_{n-1},x_{n-1}c_{n-1}x_{n}$, and
5. $c_{1}c_{2}x_{n},c_{2}c_{3}x_{n},\dots,c_{n-2}c_{n-1}x_{n}$,
then $x_{1}\rightarrow x_{n}$.
It is easy to check that the contextual transitivity rule, applied to single variables, is obtained from the above rule by setting $n=3$, $x_{1}=x$, $x_{2}=y$, $x_{3}=z$, $c_{1}=y=c_{2}$.
Proposition 22.
Proof.
We prove the case $K=\mathbb{B}$; the general case then follows from Prop. 7. Let $\mathcal{R}$ be a contextual family such that the assumptions of the contextual chain rule hold. Suppose toward a contradiction that $\mathcal{R}$ does not satisfy $x_{1}\rightarrow x_{n}$. Then $\mathcal{R}$ contains two assignments $s$ and $s^{\prime}$ in context $x_{1}c_{1}x_{n}$ such that $s(x_{1})=s^{\prime}(x_{1})$ and $s(x_{n})\neq s^{\prime}(x_{n})$. Consequently, $c_{1}\rightarrow x_{n}$ entails that $s(c_{1})\neq s^{\prime}(c_{1})$. These two assignments are illustrated in the top-left corner of Figure 5, where without loss of generality we use values 0 and $1$. By local consistency, $\mathcal{R}$ then contains two assignments $t$ and $t^{\prime}$ in the context $x_{1}c_{1}x_{2}$ such that $t(x_{1})=t^{\prime}(x_{1})$ and $t(c_{1})\neq t^{\prime}(c_{1})$. Furthermore, $x_{1}\rightarrow x_{2}$ entails $t(x_{2})=t^{\prime}(x_{2})$. These assignments are illustrated in the top-middle table of Figure 5, again using values 0 and $1$. By local consistency and induction, the remaining contexts listed in item 4 of the contextual chain rule (ccr) contain assignment pairs of the form depicted in Figure 5, with the values $b_{i}$ and $b^{\prime}_{i}$ being unknown. In particular, item 1 of ccr guarantees that the two assignments on the remaining relations agree on the variables $x_{i}$, $i\in[n]$.
Moving to the contexts listed in item 5 of ccr, the leftmost table in Figure 6 illustrates two assignments for the context $c_{1}c_{2}x_{n}$; the values of $c_{1}$ and $c_{2}$ are explained by local consistency with the top-right table of Figure 5, and the values of $x_{n}$ are explained by $c_{1}\rightarrow x_{n}$ together with local consistency with the top-left table in Figure 5. The assignments for the context $c_{2}c_{3}x_{n}$ in Figure 6 arise by local consistency with the assignments for the context $c_{2}x_{2}c_{3}$ in Figure 5, and from $c_{2}\rightarrow x_{n}$ together with local consistency with the leftmost table in Figure 6. Using this reasoning, by induction we eventually obtain the rightmost table in Figure 6.
$x_{1}$ $c_{1}$ $x_{n}$
0 0 0
0 1 1
$x_{1}$ $c_{1}$ $x_{2}$
0 0 0
0 1 0
$c_{1}$ $x_{2}$ $c_{2}$
0 0 $b_{2}$
1 0 $b^{\prime}_{2}$
$x_{2}$ $c_{2}$ $x_{3}$
0 $b_{2}$ 0
0 $b^{\prime}_{2}$ 0
$c_{2}$ $x_{2}$ $c_{3}$
$b_{2}$ 0 $b_{3}$
$b^{\prime}_{2}$ 0 $b^{\prime}_{3}$
$\dots$
$x_{n-1}$ $c_{n-1}$ $x_{n}$
0 $b_{n-1}$ 0
0 $b^{\prime}_{n-1}$ 0
Figure 5: Relations arising from Items 1, 3, and 4.
$c_{1}$ $c_{2}$ $x_{n}$
0 $b_{2}$ 0
1 $b^{\prime}_{2}$ 1
$c_{2}$ $c_{3}$ $x_{n}$
$b_{2}$ $b_{3}$ 0
$b^{\prime}_{2}$ $b^{\prime}_{3}$ 1
$\dots$
$c_{n-2}$ $c_{n-1}$ $x_{n}$
$b_{n-2}$ $b_{n-1}$ 0
$b^{\prime}_{n-2}$ $b^{\prime}_{n-1}$ 1
Figure 6: Relations arising from Items 2 and 5.
The proof of the proposition suggests that even soundness of inference rules can be non-trivial to check for general FDs.
5.2 Complexity of unary FDs and arbitrary CDs
In terms of complexity, one application of the contextual chain rule does not actually induce a higher complexity than the cycle rule from Section 3.2: we obtain $\mathrm{NL}$-completeness here as well. For the containment, note that the conditions of the contextual chain rule can be checked iteratively, by guessing the necessary $x_{i}$ and $c_{i}$ values and only keeping constantly many of them in memory at a time. For hardness, note that the first condition already requires an $x_{1}-x_{n}$ path in the underlying dependency graph. If we ensure that there exists a $c$, such that all other conditions are trivially satisfied, we can simply reduce from $\mathrm{REACH}$ again, as we did in the case for the cycle rule. The full proof can be found in the supplementary material.
Proposition 23.
The problem of deciding whether $x_{1}\rightarrow x_{n}$ can be derived from a set of CDs and unary FDs $\Sigma$ by a single application of the contextual chain rule is $\mathrm{NL}$-complete.
Even though deciding whether a single application of the chain rule yields a particular dependency can be checked in logarithmic space, the same idea does not yield a nondeterministic logspace algorithm for arbitrarily many applications of the rule. The issue is that each application of the rule yields new FDs, which need to be kept in memory, and if a derivation requires polynomially many applications of the chain rule, polynomially many FDs need to be kept in memory.
We do—however—obtain a polynomial-time algorithm by simply checking for all pairs $x,y$, if $x\rightarrow y$ is derivable from the current set of FDs and adding it, if it is. This process is repeated until no new FDs can be added. For unary FDs this only takes polynomial time.
The full proof can be found in the supplementary material.
Theorem 24.
The problem whether a unary FD $y_{1}\rightarrow y_{2}$ can be derived from a set $\Sigma$ of FDs and CDs by using the reflexivity rule, the cycle rule and the contextual chain rule is decidable in polynomial time.
6 Conclusions and further directions
We considered realisability and logical inference in a theoretical framework for modelling local consistency in the presence of (potential) global inconsistency. The framework is based on locally consistent families of $K$-relations, for positive commutative monoids $K$. We investigated when a given family $\mathcal{R}$ of locally consistent (set-based) relations can be realised as locally consistent $K$-families. In particular, we gave a complete characterisation for the case where $K$ is cancellative and $\mathcal{R}$ is a contextual family over hole context sequences. Moreover, we established a sound and complete axiomatisation for the entailment of unary FDs in the setting of locally consistent $K$-relations. We discovered two weakenings of the transitivity rule—namely, the chain and cycle rules—that are sound in the contextual setting, while the usual transitivity rule of FDs becomes unsound. Finally, we established a polynomial time algorithm for checking whether a given unary FD is derivable from a given set of CDs and unary FDs by using the reflexivity rule, the cycle rule and the contextual chain rule. For unary FDs and binary CDs, we established that the entailment problem is NL-complete. The derivability problem of unary FDs and binary CDs remains NL-hard, even if only one application of either the contextual cycle or the contextual chain rule is allowed (see Table 1 on page 1 for an overview of our complexity and axiomatisation results).
Beyond unary FDs. Until now, we have focused on the entailment problem for unary FDs. While these are the simplest kind of FDs, we have already revealed that even this case is far from trivial, and includes context specific intricacies.
We formulated our novel axiom schemas—the contextual cycle rule and the contextual chain rule—to be applied to single variables only. However, the rules remain sound even if the scopes of the rules are extended to sets of variables.
Remark 25.
The cycle rule extends from single variables to sets of variables. That is, for contextual $K$-families, if
$x_{1}\to x_{2}\to\dots\to x_{k}\to x_{1},$
then $x_{1}\to x_{k}$, for all sets of variables $x_{1},\dots,x_{k}$ with $k\geq 1$. Likewise, the contextual chain rule extends from variables to sets of variables. Specifically, the proofs of Lemma 11 and Proposition 22 do not rely on the variables being singletons, and apply verbatim to sets of variables. In particular, for a set of variables $X$, we simply interpret $t(X)=t^{\prime}(X)$ as $\forall x\in X\colon t(x)=t^{\prime}(x)$.
With the above extension, we can see that the contextual transitivity rule is an instance of the contextual chain rule by setting $n=3$, $x_{1}=x$, $x_{2}=y$, $x_{3}=z$, $c_{1}=y=c_{2}$ (as we already mentioned in Section 5). Hence reflexivity, augmentation, and the contextual chain rule constitute a sound axiomatisation in the contextual setting, that is complete for derivations within a single context (cf. Proposition 10111The proof can be adapted to the variant of CA, where the contextual transitivity rule is replaced with the contextual chain rule. The left-to-right direction will utilise the fact that Armstrong’s axioms are complete within a single context, while the right-to-left direction can be applied verbatim, since the contextual transitivity rule is an instance of the contextual chain rule.).
Future work. We conclude by listing some interesting questions deserving further study:
  • Does there exist a characterisation akin to Theorem 3 for arbitrary context sets?
  • Can a more general decomposition theorem—analogous to Proposition 18—be obtained for $\mathbb{R}_{\geq 0}$-families?
  • What is the structure of the equivalence relation on monoids that is defined by $K$-realisability of contextual families (cf. Prop. 20)? E.g., if the monoid has a non-zero idempotent element $a$ (i.e., $a+a=a$), then every contextual family is realisable. What can be said about monoids without idempotent elements, apart from 0?
  • The chain rule and the cycle rule are presented as infinite axiom schemas. Can these rules be simplified? Are they implied by some more fundamental rule? Can we replace the infinite schemas with a finite set of rules?
  • Is the full (non-unary) entailment problem for FDs invariant of the monoid (cf. Corollaries 12 and 21)? For any monoid with a non-zero idempotent element, the entailment problem coincides with that of the Boolean monoid.
  • Is the entailment problem finitely axiomatisable and decidable? For the Boolean monoid, this is a subproblem of the undecidable implication problem for FDs and inclusion dependencies (?). Note that, inclusion dependencies (see e.g., (?)) can be used to express the local consistency property (in the non-contextual setting). Hence the intricacies that arise in our axioms for FDs arise already in the non-contextual setting, if FDs are axiomatised together with inclusion dependencies.
Recent work by Hirvonen (?; ?) considers the implication problem and the axiomatisability of functional dependencies together with unary inclusion dependencies, also in the monoidal setting. As this analysis is restricted to the uni-relational setting, their results do not translate to our more general framework. Nevertheless, generalising their work to our setting may prove fruitful.
Acknowledgments
Appendix A Full proofs
@@
Proof.
The direction from right-to-left is immediate. We will prove the converse. Assume $\Sigma\vdash_{\textrm{AX}}x\rightarrow y$. Therefore there is a sequence of FDs $\psi_{0},\dots,\psi_{n-1}$ such that
  1. 1.
    $\psi_{n-1}=x\rightarrow y$
  2. 2.
    for $i\in[n]$, $\psi_{i}$ is from $\Sigma$ or obtained from $\{\psi_{0},\dots,\psi_{i-1}\}$ by using one of the rules of $\mathrm{AX}$.
  3. 3.
    for $i\in[n]$, $\mathrm{Vars}(\psi_{i})\in\textrm{Conts}(\Sigma\cup\{x\rightarrow y\})$.
For $i\in[n]$, define $\Sigma_{i}:=\{\psi_{0},\dots,\psi_{i-1}\}$. If we can show that $\Sigma_{i}\vdash_{\mathrm{CA}}\psi_{i}$, for every $i\in n$ such that $\psi_{i}\not\in\Sigma$, we can combine these derivations to obtain a proof of $\Sigma\vdash_{\mathrm{CA}}x\rightarrow y$.
To that end, let $i\in[n]$ be such that $\psi_{i}\not\in\Sigma$. Thus, $\psi_{i}$ is derivable from $\Sigma_{i}$ using some uni-context rule $(*)\in\mathrm{AX}$. Therefore, there exists a subset $\Sigma^{\prime}_{i}\subseteq\Sigma_{i}$ such that $\Sigma^{\prime}_{i}\cup\{\psi_{i}\}$ is uni-context and $\Sigma^{\prime}_{i}\vdash_{\mathrm{AX}}\psi_{i}$. Set $z:=\mathrm{Vars}(\Sigma^{\prime}_{i}\cup\{\psi_{i}\})$ and note that $z\in\textrm{Conts}(\Sigma^{\prime}_{i}\cup\{\psi_{i}\})$. Since (non-contextual) relations are special cases of contextual families and since axioms in $\mathrm{AX}$ are sound, it follows that $\Sigma^{\prime}_{i}\models\psi_{i}$ holds in the relational (non-contextual) setting. Since Armstrong’s axioms are complete, it follows that $\Sigma^{\prime}_{i}\vdash_{\mathrm{A}}\psi_{i}$. Now, by Proposition 10, we obtain $\Sigma^{\prime}_{i}\cup\{z\rightarrow z\}\vdash_{\mathrm{CA}}\psi_{i}$ and by the reflexivity axiom, that $\Sigma^{\prime}_{i}\vdash_{\mathrm{CA}}\{z\rightarrow z\}$. By Remark 9, we obtain $\Sigma^{\prime}_{i}\vdash_{\mathrm{CA}}\psi_{i}$, which implies $\Sigma_{i}\vdash_{\mathrm{CA}}\psi_{i}$. ∎
@@
Proof.
We prove the case $K=\mathbb{B}$; the proof of the general case is identical, except that all relation constructions have to be replaced with uniformly annotated $K$-relations, that is every assignment is annotated with the same monoid value in $K$. Also, note that $\mathcal{R}\models x\rightarrow y$ iff $\mathrm{Supp}(\mathcal{R})\models x\rightarrow y$ holds by definition.
First note that, if $\phi$ is a CD, then both $\Sigma\models\phi$ and $\Sigma\vdash_{\mathcal{CR}}\phi$ trivially hold. Thus we assume that $\phi$ is a unary FD. Let us consider soundness first. Suppose $(\psi_{1},\dots,\psi_{n})$ is a proof of $\psi_{n}$ from $\Sigma$, and let $\mathcal{R}$ be a contextual family over $\mathcal{C}$ such that $\mathcal{R}\models\Sigma$, and $\mathrm{Vars}(\theta)\in\mathcal{C}$ for all $\theta\in\Sigma\cup\{\psi_{n}\}$. We need to show that $\mathcal{R}\models\psi_{n}$. As the induction hypothesis, suppose $\mathcal{R}\models\psi_{i}$ for all $i<j$, where $j\leq n$. Suppose also $\psi_{j}$ is obtained from $\psi_{\ell_{1}},\dots,\psi_{\ell_{n}}$, where $\ell_{1}<\dots<\ell_{n}<j$, by using the cycle rule. In particular, the induction hypothesis entails that $\mathcal{R}$ satisfies all of $\psi_{\ell_{1}},\dots,\psi_{\ell_{n}}$ and their contexts are in $\mathcal{C}$ (due to Item 3 of Definition 8). Hence $\mathrm{Vars}(\psi_{j})\in\mathcal{C}$, due to the form of the cycle rule. From Lemma 11 it now follows that $\mathcal{R}\models\psi_{j}$, for $\mathcal{R}$ contains all the necessary contexts. The case for reflexivity is immediate by Item 3 of Definition 8, concluding the induction proof.
Case 1. There is no path from $x$ to $y$ in $G$. Let $X$ be the set of variables to which there is a path from $x$, including $x$ itself. We create one relation $R$ satisfying $\Sigma$ and falsifying $x\rightarrow y$. This relation has two assignments: $s$ that maps each variable from $V$ to 0, and $s^{\prime}$ that maps each variable from $X$ to 0 and the remaining variables from $V\setminus X$ to $1$. Since $x\in X$ and $y\in V\setminus X$, we obtain $R\not\models x\rightarrow y$. For every $u\rightarrow v\in\Sigma$, we have $R\models u\rightarrow v$, otherwise $u\in X$ and $v\in V\setminus X$, contradicting the definition of $X$. By taking the projections of $R$ on all subsets of $V$, we obtain a contextual family $\mathcal{R}$ witnessing $\Sigma\not\models x\rightarrow y$.
Case 2. There is a path from $x$ to $y$ in $G$. We construct a counterexample contextual family $\mathcal{R}$ as follows. For each $u\rightarrow v$ or $uv\rightarrow uv$ appearing in $\Sigma$ such that $\{u,v\}\neq\{x,y\}$, construct a relation that has two assignments mapping $(u,v)$ respectively to $(0,0)$ and $(1,1)$. For $x\rightarrow y$, construct a relation that has four assignments mapping $(x,y)$ respectively to $(0,0),(0,1),(1,0),(1,1)$. The local consistency criterion is satisfied, since each variable has values 0 and $1$ in every relation in which it appears. Let $\mathcal{R}$ denote the contextual family that arises from the aforementioned relations. Clearly $\mathcal{R}\not\models x\rightarrow y$, but $\mathcal{R}\models u\rightarrow v$, for each $u\rightarrow v\in\Sigma$ such that $\{u,v\}\neq\{x,y\}$. It remains to consider the case of $u\rightarrow v\in\Sigma$, where $\{u,v\}=\{x,y\}$. Clearly, it cannot be that $u\rightarrow v=x\rightarrow y$, since this would contradict the assumption that $\Sigma\not\vdash x\rightarrow y$. Moreover, if $u\rightarrow v=y\rightarrow x$, then the path from $x$ to $y$ and the cycle rule entails $\Sigma\vdash x\rightarrow y$, contradicting our assumption. Hence we conclude that all FDs in $\Sigma$ are satisfied. The resulting contextual family $\mathcal{R}$ thus provides a proof for $\Sigma\not\models x\rightarrow y$, concluding the second case and the proof. In the general case of $K\neq\mathbb{B}$, using positivity and non-triviality of $K$ we can select two non-zero monoid values $a,b\in K$ such that $a+a=b$. Then assignments in the above defined ($K$-)relations of cardinality $2$ shall be annotated with $b$, while assignments in the ($K$-)relation of cardinality $4$ shall be annotated with $a$. It is easy to see that with this amendment the proof goes through also in the general case. ∎
For the proof of Theorem 3, we need some auxiliary concepts and results.
A digraph $G=(V,E)$ is called cyclically $n$-partite if there exists a partition $\{V_{0},\dots,V_{n-1}\}$ of $V$ such that for every edge $(u,v)\in E$ we have $u\in V_{i}$ and $v\in V_{i+1}$, for some $i\in[n-1]$. The following lemma is a simple observation.
@@
Let $\mathcal{R}$ be a contextual family over a hole $\mathcal{C}=(C_{0},\dots,C_{n-1})$. Then $\mathrm{OPGraph}(\mathcal{R},\mathcal{C})$ is cyclically $n$-partite.
Proof.
The claim follows from the definition of $\mathrm{OPGraph}(\mathcal{R},\mathcal{C})$, and the required partition is
$\{\mathcal{R}_{1}\!\restriction\!(C_{1}\cap C_{2}),\dots,\mathcal{R}_{n-1}\!\restriction\!(C_{n-1}\cap C_{n}),\mathcal{R}_{n}\!\restriction\!(C_{n}\cap C_{1})\}.$
Then, we show that every cycle gives rise to a uniformly weighted $K$-relation. Recall that, by definition, a (simple) cycle $(v_{1},\dots,v_{\ell},v_{1})$ does not visit any vertex twice, except for the first and the last one. We say that a contextual family $\mathcal{R}$ over a hole is cyclic if $\mathrm{OPGraph}(\mathcal{R},\mathcal{C})$ contains a cycle
@@
where $\bigcup{\mathcal{R}}=\{s_{1},\dots,s_{\ell}\}$. The next lemma states that any cyclic family can be realised as a $K$-family using uniform multiplicities.
Lemma 27.
Let $\mathcal{R}$ be a cyclic contextual family over a hole $\mathcal{C}=(C_{0},\dots,C_{n-1})$. Let $K$ be a positive commutative monoid and $0\neq w\in K$. Then $w\mathcal{R}$ is a contextual $K$-family, and $\mathrm{Supp}(w\mathcal{R})=\mathcal{R}$.
@@
We need to prove that $w\mathcal{R}$ is locally consistent. Suppose the cycle has length $\ell$, as in (1). Since $\mathrm{OPGraph}(\mathcal{R},\mathcal{C})$ is cyclically $n$-partite (Lemma 26), $\ell$ must be a multiple of $n$; that is, $n\cdot m=\ell$ for some positive integer $m$. Fix $i\in[n]$. One observes that
@@
[…]
@@
( $\Rightarrow$ ) Let $\mathcal{R}$ be the support of some contextual $K$-family $\mathcal{S}$. Then $\mathcal{S}$ can be represented as $\colon\bigcup{\mathcal{R}}\to K_{\neq 0}$, a mapping which we extend to sets $S$ by $(S)\coloneqq\sum_{s\in S}\#(s)$. Toward a contradiction, assume $\mathrm{OPGraph}(\mathcal{R},\mathcal{C})$ is not edge-cyclic. Then, some $s\in\bigcup{\mathcal{R}}$ gives rise to an edge that is not in any cycle. Without loss of generality, we assume that $s$ is associated with the context $C_{0}$, that is, it belongs to $\mathcal{R}_{0}$. We define sets of assignments $L^{j}$, $j\geq 0$, inductively as follows:
  • $L^{0}\coloneqq\{s^{\prime}\in\mathcal{R}_{{1}}\mid{s}\!\restriction\!{(C_{0}\cap C_{1})}={s^{\prime}}\!\restriction\!{(C_{0}\cap C_{1})}\}$,
  • $L^{j+1}$ extends $L^{j}$ with assignments $s^{\prime}\in\mathcal{R}_{{j+1}}$, $j\in[n]$, such that there is $s\in L^{i}\cap\mathcal{R}_{{j}}$ with
@@
$\hskip-5.69054pt\#(M_{0}\cup\{s\})\leq\#(M_{1})\leq\ldots\leq\#(M_{n-1})\leq\#(M_{0}),$ (2)
where $\leq$ refers to the natural (pre)order of $K$. Note that if this claim holds, we reach a contradiction. Indeed, note that $s\notin L_{0}$ by hypothesis, hence the transitivity of $\leq$ implies $(M_{0})+\#(\{s\})=\#(M_{0}\cup\{s\})\leq\#(M_{0})$. Thus, there is $c\in K$ such that $(M_{0})+\#(\{s\})+c=\#(M_{0})$. By cancellativity, $(\{s\})+c=0$, and hence by positivity, #(s)=0\#(s)=0. This contradicts the fact that $s\in\bigcup{\mathcal{R}}$, where #\# is a function from $\bigcup{\mathcal{R}}$ to $K_{\neq 0}$.